# Review: Quantum Metrology and Sensing with Many-Body Systems

### v1

@@ V Equilibrium Quantum Sensing: Localization Transition
So far, second-order quantum phase transition in the ground state of many-body systems, manifested in different setups, has been investigated as a resource of enhanced precision sensing. In contrast, localization transitions, exemplified through Anderson and many-body localizations, affect the entire spectrum of the system and thus their impact is more drastic. Originally disorder is used for suppressing the effect of particle tunneling and thus inducing localization in which particles’ wave functions extends only locally across a few sites. Localization takes place in both non-interacting (Anderson localization) and interacting (many-body localization) systems. The localization transition is energy dependent such that each eigenstate localizes at a different value of disorder Luitz et al. (2015); Guo et al. (2021a); Yousefjani and Bayat (2023). Localization transition can also be induced by Floquet dynamics Yousefjani et al. (2023b) pseudo-random potentials Zhang and Yao (2018); Žnidarič and Ljubotina (2018) or Stark fields Schulz et al. (2019); Morong et al. (2021). In this section, we investigate the application of both single- and multi-particle systems under the effect of such localization effects for quantum sensing.
@@ V.2 Stark localization
For one particle that tunnels to its neighbors in a one-dimensional lattice affected by a linear gradient field with strength $h$, the Hamiltonian is given by
@@
where $J$ is the tunneling rate. In this setup, the gradient-field strength $h$ is the unknown parameter and aimed to be estimated. It is well known that in the thermodynamic limit (i.e. $L{\rightarrow}\infty$), all energy levels of the Hamiltonian Eq.(56) go through a phase transition from an extended to a localized phase at $h_{c}{=}0$ van Nieuwenburg et al. (2019); Kolovsky (2008); Schulz et al. (2019); Chanda et al. (2020); Yao and Zakrzewski (2020). Focusing on the ground state, Fig. 5 (a) presents QFI as a function of $h$ for different sizes of the probe. The initial plateaus represent the extended phase in which the QFI takes its maximum for $h{\leq}h_{\max}$. Clearly, $h_{\max}$ skews to zero and tends to the critical point $h_{c}{=}0$, by increasing the length of the probe. The maximum values of the QFI, denoted by $I_{Q}(h_{\max})$ increase dramatically by enlarging the probe, evidencing divergence of the QFI in the thermodynamic limit. In the localized phase, namely $h{>}h_{\max}$, the QFI is size-independent and decays algebraically according to $I_{Q}{\propto}|h-h_{\max}|^{-\alpha}$ with $\alpha{=}2$ for the ground state. The scaling of the QFI with the system size at the critical point has been plotted in Fig. 5 (b). The QFI shows strong super-Heisenberg scaling, i.e. $I_{Q}(h_{\max}){\propto}L^{\beta}$ with $\beta{\simeq}6$. The inset of Fig. 5 (b) shows how $\beta$ changes by getting distance from the critical point, i.e. $h_{\max}$. As the Stark localization transition happens in the entire energy spectrum of Eq. (56), one can inquire about higher energy eigenstates’ scaling. Interestingly, one can obtain the super-Heisenberg scaling with $\beta{\simeq}4$ for all excited eigenstates. Following section IV, establishing finite-size scaling analysis results in critical exponents $(\alpha,\nu){=}(2.00,0.33)$ and $(\alpha,\nu){=}(4.00,1.00)$ for the ground state and a typical eigenstate from the mid-spectrum, respectively. Indeed, three exponents $\beta,\alpha$ and $\nu$ obey $\beta{=}\alpha/\nu$ in Eq. (42), regardless of the energy level.
@@ VI.2 Non-Hermitian Quantum Sensors
where $E_{n}$ is the corresponding eigenvalue. However, these states are biorthogonal Brody (2013), and upon normalization gives, $\braket{\tilde{\psi}^{\rm L}_{m}}{\tilde{\psi}^{\rm R}_{n}}{=}\delta_{mn}$. Here, |$\ket{\tilde{\psi}^{\rm R}_{n}}{=}\ket{\psi^{\rm R}_{n}}/\sqrt{\braket{\psi^{\rm L}_{n}}{\psi^{\rm R}_{n}}}$, and |$\ket{\tilde{\psi}^{\rm L}_{n}}{=}\ket{\psi^{\rm L}_{n}}/\sqrt{\braket{\psi^{\rm L}_{n}}{\psi^{\rm R}_{n}}}$. This helps in writing down the completeness relation $\sum_{n}\ket{\tilde{\psi}^{\rm R}_{n}}\bra{\tilde{\psi}^{\rm L}_{n}}{=}\mathds{1}$ and the spectral decomposition $H_{\rm NH}{=}\sum_{n}E_{n}\ket{\tilde{\psi}^{\rm R}_{n}}\bra{\tilde{\psi}^{\rm L}_{n}}$ |. Non-Hermitian Hamiltonians can have degeneracies analogous to Hermitian systems, i.e. identical eigenvalues with distinct eigenstates. The points in the Hamiltonian’s parameter space where this happens are known as diabolic points. There are also parameter values where more than one eigenvalues and the corresponding eigenstates can become identical. These are known as Exceptional Points (EP) and have no Hermitian counterpart. The right- and left-eigenstates corresponding to the EP are orthogonal to each other and the biorthogonal framework breaks down as the Hamiltonian is not digonalizable anymore. If the Hamiltonian has parity-time (PT) symmetry, then in certain parameter regime, the eigenvalues are real and the Hamiltonian and the PT operator share the same eigenstates. This is known as the PT-unbroken phase. In the PT-broken phase, the eigenvalues can be complex. The PT-symmetry breaking point is an EP of the system.
@@
EPs are branch point singularities in the complex energy manifold and generate non-trivial topology, quantified by winding of eigenvalues and eigenvectors Leykam et al. (2017). Sensing based on EP was proposed first in Ref. Wiersig (2014) for coupled cavity systems – with photon loss and gain terms – that have resonant modes at EP. A non-Hermitian perturbation with strength $\epsilon$ would move the system away from the EP. For an EP of degree $n$ (i.e. $n$ eigenstates coalescing), the resulting energy splitting ${\sim}\epsilon^{1/n}$. In contrast, for Hermitian degeneracies or diabolic points, the leading order splitting is ${\sim}\epsilon$. The advantage of the EP based sensors can be quantified by the susceptibility $d_{\epsilon}\epsilon^{1/n}=\epsilon^{-(1-1/n)}$, which diverges as $\epsilon\to 0$. This was experimentally demonstrated with optical microring resonators Chen et al. (2017); Hodaei et al. (2017) and also with further imposition of PT symmetry Liu et al. (2016); Yu et al. (2020). However, these schemes do not reveal the precision of the sensor which is the smallest change in $\epsilon$ that can be measured. The standard measurement is done by first implementing an excitation, followed by analyzing the resonant scattering output. The signal-to-noise ratio of the output is a measure of QFI, which sets the precision. As the eigenstates participating in the excitation-deexcitation process become more and more non-orthogonal while approaching EP, it was debated that the resulting quantum noise would overwhelm the advantage of EP based sensors Langbein (2018); Chen et al. (2019). Experimental observation also confirmed this Wang et al. (2020). Signal enhancement was nevertheless predicted at specific frequencies Zhang et al. (2019) or with additional gain terms Lau and Clerk (2018) and was also observed experimentally for systems slightly detuned from EP Kononchuk et al. (2022). The effectiveness of EP based sensors is still a topic of active research Wiersig (2020).
Figure 9: Non-Hermitian energy gaps. Two types of gaps can occur in the complex energy plane (middle panel). Line gap (blue dashed line) separates two bands (red loops). Point gap resides inside a spectral loop (e.g. $E_{0}$). Line gap closes when the loops merge (left panel). Point gap closes when the loop shrinks to an arc (right panel).
@@
Beyond sensing boundary perturbations, non-Hermitian topology also can be used for estimating bulk Hamiltonian parameters. The notion of band gaps are extended in the non-Hermitian regime to two different gap structures: line gap and point gap (see Fig. 9Kawabata et al. (2019). A spectral loop (for 1D systems) or finite spectral area (in higher dimensions) correspond to the presence of non-Hermitian skin effect in finite systems, where most of the eigenstates are localized at a boundary Okuma et al. (2020); Zhang et al. (2020, 2022); Li et al. (2020). Changes in the bulk Hamiltonian parameters can make the spectral structure contract to a curve with zero spectral area. This signals a closure of the point gap, which makes the skin effect vanish. The corresponding localization transition in the eigenstates can be used for sensing, as was shown in Ref. Sarkar et al. (2024), to achieve precision with Heisenberg scaling. This extends the role of gap closing for quantum-enhanced sensitivity to a truly non-Hermitian domain. For the aforementioned Hatano-Nelson model, this scaling can be obtained analytically (see Fig. 10). The robustness against local disorders is also shown by mapping it to a non-Hermitian Aubrey-André-Harper model Longhi (2019). Apart form extension to prototypical multi-band models Yao and Wang (2018) and 2D systems Yao et al. (2018), experimental implementation is also possible in discrete-time non-Unitary quantum walk setups Xiao et al. (2020, 2021).
@@ VII Resource analysis for critical quantum sensing
As discussed above, various forms of criticality can achieve quantum enhanced sensitivity. The resource for achieving such enhanced precision is the system size $N$ and the figure of merit is the QFI. However, in some cases the time $T$ needed for accomplishing the sensing task has to also be considered as a resource. This might be due to practical issues, such as limited coherence time, or size dependence $T$ which makes the sensing time extremely long for large system sizes. To account for such complexity, one can use normalized QFI as a figure of merit, namely $I_{Q}/T$ Albarelli et al. (2018); Rossi et al. (2020); Chin et al. (2012); Chaves et al. (2013); Brask et al. (2015); Smirne et al. (2016). For instance, in the case of many-body ground state sensing, one may consider an adiabatic approach for generating the complex ground state. This can be achieved by slowly evolving the system initialized in the ground state of a simpler Hamiltonian which is transformed into the desired Hamiltonian over the time interval $T$. If the transformation is slow enough, namely $T$ is large, the quantum state remains in the ground state of the instant Hamiltonian throughout the evolution. This guarantees that at the end of evolution when the Hamiltonian takes the final form the system is in the desired ground state. When the system is adiabatically driven towards a second-order quantum phase transition, the required time $T$ scales as $T{\sim}N^{z}$ Rams et al. (2018), where $z$ is the dynamical critical exponent showing how the energy gap $\Delta E$ vanishes at the critical point by increasing the system size $\Delta E_{\rm min}\sim 1/N^{z}$. Incorporating this into the scaling of the QFI across second-order quantum phase transitions in Eq. (39) leads to
@@
Similar results can be found for different quantum sensing strategies where the required time depends on the system size. For instance, in Ref. Montenegro et al. (2023) quantum enhanced sensitivity, i.e. $\beta{>}1$, is achieved for the steady state at the boundary time crystal transition when the figure of merit is the QFI $I_{Q}$. However, if one changes their figure of merit to normalized QFI, i.e. $I_{Q}/T$, with $T$ being the time that the system reaches its steady state, the scaling with respect to system size is reduced to standard limit. However this situation changes for Stark probe presented in section V.2. In Ref  He et al. (2023); Yousefjani et al. (2023a), it has been shown that regardless of the number of particles and the range of the interaction, one always has $\beta-z{>}1$ ensuring that the quantum enhanced sensitivity is always achievable for the Stark probes.
Note that a true resource analysis depends on the sensing scheme. If the time is not a restricting issue QFI can be used as the figure of merit. Equivalently, for those sensing schemes that time is a key limiting factor a better figure of merit is normalized QFI, namely $I_{Q}/T$.
@@ VIII Non-Equilibrium Quantum Sensing: Dynamical Quantum Phase Transitions
The footprints of the ground state quantum criticality in equilibrium can also be observed in non-equilibrium dynamics. For a closed system described by a time-independent Hamiltonian $H$, its time evolution which is described by Schrödnger equation as

### v2

@@ V Equilibrium Quantum Sensing: Localization Transition
So far, second-order quantum phase transition in the ground state of many-body systems, manifested in different setups, has been investigated as a resource of enhanced precision sensing. In contrast, localization transitions, exemplified through Anderson and many-body localizations, affect the entire spectrum of the system and thus their impact is more drastic. Originally disorder was used for suppressing the effect of particle tunneling and thus inducing localization in which particles’ wave functions extend only locally across a few sites. Localization takes place in both non-interacting (Anderson localization) and interacting (many-body localization) systems. The many-body localization transition is energy dependent such that each eigenstate localizes at a different value of disorder Luitz et al. (2015); Guo et al. (2021a); Yousefjani and Bayat (2023). Localization transition can also be induced by Floquet dynamics Yousefjani et al. (2023b) pseudo-random potentials Zhang and Yao (2018); Žnidarič and Ljubotina (2018) or Stark fields Schulz et al. (2019); Morong et al. (2021). In this section, we investigate the application of both single- and multi-particle systems under the effect of such localization effects for quantum sensing.
@@ V.2 Stark localization
For a single particle that tunnels to its neighbors in a one-dimensional lattice affected by a linear gradient field with strength $h$, the Hamiltonian is given by
@@
where $J$ is the tunneling rate. In this setup, the gradient-field strength $h$ is the unknown parameter and aimed to be estimated. It is well known that in the thermodynamic limit (i.e. $L{\rightarrow}\infty$), all energy levels of the Hamiltonian Eq.(56) go through a phase transition from an extended to a localized phase at $h_{c}{=}0$ van Nieuwenburg et al. (2019); Kolovsky (2008); Schulz et al. (2019); Chanda et al. (2020); Yao and Zakrzewski (2020). Focusing on the ground state, Fig. 5 (a) presents QFI as a function of $h$ for different sizes of the probe. The initial plateaus represent the extended phase in which the QFI takes its maximum for $h{\leq}h_{\max}$. Clearly, $h_{\max}$ skews to zero and tends to the critical point $h_{c}{=}0$, as the length of the probe increases. The maximum values of the QFI, denoted by $I_{Q}(h_{\max})$, increase dramatically by enlarging the probe, evidencing divergence of the QFI in the thermodynamic limit. In the localized phase, namely $h{>}h_{\max}$, the QFI is size-independent and decays algebraically according to $I_{Q}{\propto}|h-h_{\max}|^{-\alpha}$ with $\alpha{=}2$ for the ground state. The scaling of the QFI with the system size at the critical point has been plotted in Fig. 5 (b). The QFI shows strong super-Heisenberg scaling, i.e. $I_{Q}(h_{\max}){\propto}L^{\beta}$ with $\beta{\simeq}6$. The inset of Fig. 5 (b) shows how $\beta$ changes by getting distance from the critical point, i.e. $h_{\max}$. As the Stark localization transition happens in the entire energy spectrum of Eq. (56), one can inquire about higher energy eigenstates’ scaling. Interestingly, one can obtain the super-Heisenberg scaling with $\beta{\simeq}4$ for all excited eigenstates. Following section IV, establishing finite-size scaling analysis results in critical exponents $(\alpha,\nu){=}(2.00,0.33)$ and $(\alpha,\nu){=}(4.00,1.00)$ for the ground state and a typical eigenstate from the mid-spectrum, respectively. Indeed, three exponents $\beta,\alpha$ and $\nu$ obey $\beta{=}\alpha/\nu$ in Eq. (42), regardless of the energy level.
@@ VI.2 Non-Hermitian Quantum Sensors
where $E_{n}$ is the corresponding eigenvalue. However, these states are biorthogonal Brody (2013), and upon normalization gives, $\braket{\tilde{\psi}^{\rm L}_{m}}{\tilde{\psi}^{\rm R}_{n}}{=}\delta_{mn}$. Here, |$\ket{\tilde{\psi}^{\rm R}_{n}}{=}\ket{\psi^{\rm R}_{n}}/\sqrt{\braket{\psi^{\rm L}_{n}}{\psi^{\rm R}_{n}}}$, and |$\ket{\tilde{\psi}^{\rm L}_{n}}{=}\ket{\psi^{\rm L}_{n}}/\sqrt{\braket{\psi^{\rm L}_{n}}{\psi^{\rm R}_{n}}}$. This helps in writing down the completeness relation $\sum_{n}\ket{\tilde{\psi}^{\rm R}_{n}}\bra{\tilde{\psi}^{\rm L}_{n}}{=}\mathds{1}$ and the spectral decomposition $H_{\rm NH}{=}\sum_{n}E_{n}\ket{\tilde{\psi}^{\rm R}_{n}}\bra{\tilde{\psi}^{\rm L}_{n}}$ |. Non-Hermitian Hamiltonians can have degeneracies analogous to Hermitian systems, i.e. identical eigenvalues with distinct eigenstates. The points in the Hamiltonian’s parameter space where this happens are known as diabolic points. There are also parameter values where more than one eigenvalues and the corresponding eigenstates can become identical. These are known as Exceptional Points (EP) and have no Hermitian counterpart. The right- and left-eigenstates corresponding to the EP are orthogonal to each other and the biorthogonal framework breaks down as the Hamiltonian is not digonalizable anymore. If the Hamiltonian has parity-time (PT) symmetry, then it commutes with the antilinear and antiunitary $\mathcal{PT}$ operator. For example, in spin-$1{/}2$ systems on a lattice, the parity operator is $\mathcal{P}{=}\sum_{j}\ket{-j}\bra{j}\otimes\sigma^{x}_{j}$, and the time-reversal operator is $\mathcal{T}{=}\sum_{j}\ket{j}\bra{j}\otimes\sigma^{y}_{j}K_{j}$. Here the term $|{-j}\rangle\langle j|$ swaps the position of spins at sites $j$ and $-j$. The action of $K_{j}$ is complex conjugation, and $\sigma^{(x,y,z)}_{j}$ are the Pauli spin operators on $j$-th site that affect the spin degree of freedom Krasnok et al. (2021); Mochizuki et al. (2016). In an appropriate parameter regime, a PT-symmetric Hamiltonian supports a real energy spectrum, and it is known as the PT-unbroken phase. In this case, the Hamiltonian and the PT operator share the same eigenstates, known as PT-symmetric eigenstates. Upon variation of the Hamiltonian parameters, a PT-broken phase can be entered where some or all eigenvalues become complex that appear in complex conjugate pairs. This happens through a phase transition associated with spontaneous PT-symmetry breaking. The PT-symmetry breaking point is an EP of the system.
@@
EPs are branch point singularities in the complex energy manifold and generate non-trivial topology, quantified by winding of eigenvalues and eigenvectors Leykam et al. (2017). Sensing based on EP was proposed first in Ref. Wiersig (2014) for coupled cavity systems – with photon loss and gain terms – that have resonant modes at EP. A non-Hermitian perturbation with strength $\epsilon$ would move the system away from the EP. For an EP of degree $n$ (i.e. $n$ eigenstates coalescing), the resulting energy splitting ${\sim}\epsilon^{1/n}$. In contrast, for Hermitian degeneracies or diabolic points, the leading order splitting is ${\sim}\epsilon$. The advantage of the EP based sensors can be quantified by the susceptibility $d_{\epsilon}\epsilon^{1/n}\propto\epsilon^{-(1-1/n)}$, which diverges as $\epsilon\to 0$. This was experimentally demonstrated with optical microring resonators Chen et al. (2017); Hodaei et al. (2017) and also with further imposition of PT symmetry Liu et al. (2016); Yu et al. (2020). However, these schemes do not reveal the precision of the sensor which is the smallest change in $\epsilon$ that can be measured. The standard measurement is done by first implementing an excitation, followed by analyzing the resonant scattering output. The signal-to-noise ratio of the output is a measure of QFI, which sets the precision. As the eigenstates participating in the excitation-deexcitation process become more and more non-orthogonal while approaching EP, it was debated that the resulting quantum noise would overwhelm the advantage of EP based sensors Langbein (2018); Chen et al. (2019). Experimental observation also confirmed this Wang et al. (2020). Signal enhancement was nevertheless predicted at specific frequencies Zhang et al. (2019) or with additional gain terms Lau and Clerk (2018) and was also observed experimentally for systems slightly detuned from EP Kononchuk et al. (2022). The effectiveness of EP based sensors is still a topic of active research Wiersig (2020).
Figure 9: Non-Hermitian energy gaps. Two types of gaps can occur in the complex energy plane (middle panel). Line gap (blue dashed line) separates two bands (red loops). Point gap resides inside a spectral loop (e.g. $E_{0}$). Line gap closes when the loops merge (left panel). Point gap closes when the loop shrinks to an arc (right panel). Figure taken from Ref. Sarkar et al. (2024).
@@
Beyond sensing boundary perturbations, non-Hermitian topology also can be used for estimating bulk Hamiltonian parameters. The notion of band gaps are extended in the non-Hermitian regime to two different gap structures: line gap and point gap (see Fig. 9Kawabata et al. (2019). A spectral loop (for 1D systems) or finite spectral area (in higher dimensions) correspond to the presence of non-Hermitian skin effect in finite systems Martinez Alvarez et al. (2018); Yao and Wang (2018), where most of the eigenstates are localized at a boundary Borgnia et al. (2020); Zhang et al. (2020, 2022); Li et al. (2020). Changes in the bulk Hamiltonian parameters can make the spectral structure contract to a curve with zero spectral area. This signals a closure of the point gap, which makes the skin effect vanish. The corresponding localization transition in the eigenstates can be used for sensing, as was shown in Ref. Sarkar et al. (2024), to achieve precision with Heisenberg scaling. This extends the role of gap closing for quantum-enhanced sensitivity to a truly non-Hermitian domain. For the aforementioned Hatano-Nelson model, this scaling can be obtained analytically (see Fig. 10). The robustness against local disorders is also shown by mapping it to a non-Hermitian Aubrey-André-Harper model Longhi (2019). Apart form extension to prototypical multi-band models Yao and Wang (2018) and 2D systems Yao et al. (2018), experimental implementation is also possible in discrete-time non-Unitary quantum walk setups Xiao et al. (2020, 2021).
@@ VII Resource analysis for critical quantum sensing
As discussed above, various forms of criticality can achieve quantum enhanced sensitivity. The resource for achieving such enhanced precision is the system size $N$ and the figure of merit is the QFI. However, in some cases the time $T$ needed for accomplishing the sensing task also has to be considered as a resource. This might be due to practical issues, such as limited coherence time, or size dependent $T$ which makes the sensing time extremely long for large system sizes. To account for such complexity, one can use normalized QFI as a figure of merit, namely $I_{Q}/T$ Albarelli et al. (2018); Rossi et al. (2020); Chin et al. (2012); Chaves et al. (2013); Brask et al. (2015); Smirne et al. (2016). For instance, in the case of many-body ground state sensing, one may consider an adiabatic approach for generating the complex ground state. This can be achieved by slowly evolving the system initialized in the ground state of a simpler Hamiltonian which is transformed into the desired Hamiltonian over the time interval $T$. If the transformation is slow enough, namely $T$ is large, the quantum state remains in the ground state of the instant Hamiltonian throughout the evolution. This guarantees that at the end of evolution when the Hamiltonian takes the final form the system is in the desired ground state. When the system is adiabatically driven towards a second-order quantum phase transition, the required time $T$ scales as $T{\sim}N^{z}$ Rams et al. (2018), where $z$ is the dynamical critical exponent showing how the energy gap $\Delta E$ vanishes at the critical point by increasing the system size $\Delta E_{\rm min}\sim 1/N^{z}$. Incorporating this into the scaling of the QFI across second-order quantum phase transitions in Eq. (39) leads to
@@
Similar results can be found for different quantum sensing strategies where the required time depends on the system size. For instance, in Ref. Montenegro et al. (2023) quantum enhanced sensitivity, i.e. $\beta{>}1$, is achieved for the steady state at the boundary time crystal transition when the figure of merit is the QFI $I_{Q}$. However, if one changes their figure of merit to normalized QFI, i.e. $I_{Q}/T$, with $T$ being the time that the system reaches its steady state, the scaling with respect to system size is reduced to standard limit.
The time needed for accomplishing the sensing task, is directly related to the strategy that one exploits. In Ref. Gietka et al. (2021), the authors show that by adiabatic evolution, one cannot get advantage for quantum sensing. However, the proof relies on a specific choice of the adiabatic evolution. In their formalism the Hamiltonian is written as $H{=}\theta H_{0}{+}H_{1}(t)$, where $H_{0}$ and $H_{1}$ do not depend on the target parameter $\theta$. In this scenario, $H_{0}$ is always on and the time dependent Hamiltonian $H(t)$ varies slowly. The resulted dynamics encodes the information about $\theta$ into the state of the system whose initial state does not depend on $\theta$. In this case, as proved in Ref. Gietka et al. (2021), there is no quantum advantage. Nonetheless, one may consider to prepare the probe in the ground state of $H_{1}$ and as $H_{1}$ is turned off, the probe is put under the action of a general $H_{0}(\theta)$ adiabatically to prepare the probe in the ground state of a general $H_{0}(\theta)$. In this situation, obtaining quantum advantage is possible. For instance, in the case of Stark systems, presented in section V.2, it has been shown that regardless of the number of particles and the range of the interaction, one always has $\beta-z{>}1$ ensuring that the quantum enhanced sensitivity is always achievable for the Stark probes He et al. (2023); Yousefjani et al. (2023a).
Note that a true resource analysis depends on the sensing scheme. If the time is not a restricting issue QFI can be used as the figure of merit. Equivalently, for those sensing schemes where time is a key limiting factor, a better figure of merit is normalized QFI, namely $I_{Q}/T$.
@@ VIII Non-Equilibrium Quantum Sensing: Dynamical Quantum Phase Transitions
The footprints of the ground state quantum criticality in equilibrium can also be observed in non-equilibrium dynamics. For a closed system described by a time-independent Hamiltonian $H$, its time evolution which is described by Schrödinger equation as